OpenStudy (anonymous):

what is the general solution for sinx+cosx = 1

OpenStudy (sidsiddhartha):

divide both sides by\[\frac{ 1 }{ \sqrt{2}}\]

OpenStudy (anonymous):

aliter sinx - (1-cosx) = 0 2sin(x/2)cos(x/2) -2(sin(x/2))^2 =0 2sin(x/2) * (sin(x/2) -cos(x/2)) =0 thus sin(x/2) = 0 or (sin(x/2) -cos(x/2)) =0 ==> tan(x/2) =1 hence further..

OpenStudy (anonymous):

if sin(x/2) = 0 then x/2 = n(pi) where n is an integer... or x = 2npi

OpenStudy (anonymous):

if tan(x/2) =1 = tan(pi/4) then x = mpi +(pi/4) m is any integer